Alpha Pharmaceuticals has a patent on a new medication used to treat high blood pressure, so it is the monopoly seller of this new drug product. The marginal cost of producing one dose of the drug is R10, and the elasticity of demand for the product is -3. What is the profit maximising monopoly price for this patented drug product


1
Expert's answer
2022-03-21T16:09:51-0400

In this case we use the formula;

"MR= MC"

"MR= P(1- \frac{1}{|e|}) = MC"

"10= P+\frac{P}{3}"

"30=3P+P"

"30=4P"

"P = 7.5"

The monopoly price for the patented drug is R7.5